OpenStudy (anonymous):

What is x^6/x^? = x^4? What is the x^? equal?

OpenStudy (anonymous):

When you divide numbers wit the same base x in this case, you subtract the power. x^6/x^2 = x^4. 6-2=4

OpenStudy (anonymous):

Thanks!